# Privacy of the last iterate in cyclically-sampled DP-SGD on nonconvex composite losses

### v1

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) refers to a family of optimization algorithms that provide a guaranteed level of differential privacy (DP) through DP accounting techniques. However, current accounting techniques make assumptions that diverge significantly from practical DP-SGD implementations. For example, they may assume the loss function is Lipschitz continuous and convex, sample the batches randomly with replacement, or omit the gradient clipping step.
In this work, we analyze the most commonly used variant of DP-SGD, in which we sample batches cyclically with replacement, perform gradient clipping, and only release the last DP-SGD iterate. More specifically — without assuming convexity, smoothness, or Lipschitz continuity of the loss function — we establish new Rényi differential privacy (RDP) bounds for the last DP-SGD iterate under the mild assumption that (i) the DP-SGD stepsize is small relative to the topological constants in the loss function, and (ii) the loss function is weakly-convex. Moreover, we show that our bounds converge to previously established convex bounds when the weak-convexity parameter of the objective function approaches zero. In the case of non-Lipschitz smooth loss functions, we provide a weaker bound that scales well in terms of the number of DP-SGD iterations.
@@ 1 Introduction
Tight privacy guarantees for DP-SGD are challenging to achieve in real-world implementations because existing approaches rely on simplifying assumptions. Some common ones include convexity and Lipschitz continuity of the loss function, randomly sampling the batches, and omitting the ubiquitous gradient clipping step. More crucially, existing practical methods assume that the weights of a trained model are always released at every DP-SGD iteration.
In this work, we develop a family of RDP bounds on the last iterate of DP-SGD. Moreover, these bounds are novel in that they:
@@
    are developed under a DP-SGD gradient clipping step without assuming that the objective function is Lipschitz continuous;
@@
    require substantially weaker assumptions compared to existing works;
@@
    are parameterized by a weak convexity parameter $m\geq 0$, for which the bounds smoothly converge to similar ones in the convex case as $m\to 0$ when the loss function has a Lipschitz-smooth component; and
  • (iv)
    can be made smaller by decreasing the SGD stepsize, increasing the standard deviation in the DP-SGD Gaussian noise, or increasing the batch size.
Problem definition. More precisely, we develop RDP bounds for the last iterates of DP-SGD applied to the composite optimization problem
@@
where $h$ is convex and proper lower-semicontinuous and $f_{i}$ is continuously differentiable on the domain of $h$. Notice that the assumption on $h$ encapsulates (i) common nonsmooth regularization functions such as the $\ell_{1}$-norm $\|\cdot\|_{1}$, nuclear matrix norm $\|\cdot\|_{*}$, and elastic net regularizer and (ii) indicator functions on possibly unbounded closed convex sets. A common setting in the wild is when $(1/k)\sum_{i=1}^{k}f_{i}(x)$ corresponds to a softmax cross-entropy loss function and $h(x)$ corresponds to an $\ell_{1}$- or $\ell_{2}$-regularization function.
Types of bounds. Each of the established bounds is of the form
$D_{\alpha}(X_{T}\|X_{T}^{\prime})\preceq\frac{\alpha\cdot{\cal{B}}_{\lambda}(C,b,T,\ell)}{\sigma^{2}}$ (2)
for some finite function $\cal B_{\lambda}(\cdot)$ and $\lambda>0$, where (i) $D_{\alpha}(X\|X^{\prime})$ denotes the $\alpha$-Rényi divergence between random variables $X$ and $X^{\prime}$, (ii) $X_{T}$ and $X_{T}^{\prime}$ are the last DP-SGD iterates on neighboring datasets, (iii) and the paramaters $\sigma$, $\lambda$, $C$, $b$, $T$, and $\ell$ are DP-SGD’s standard deviation, stepsize, $\ell_{2}$ clip norm, batch size, number of steps, and number of steps per dataset pass, respectively.
More specifically, we target three regimes. First, without any additional assumptions, we obtain (2) with $B_{\lambda}(C,b,T,\ell)=T(\lambda C)^{2}$. Second, assuming that the DP-SGD iterates are contained within an $\ell_{2}$ ball of diameter $d_{h}$, we obtain (2) with $B_{\lambda}(C,b,T,\ell)=(d_{h}+\lambda C/b)^{2}$ for small enough $\lambda$. Third, assuming that each $\nabla f_{i}$ is Lipschitz continuous, we obtain (2) with
$B_{\lambda}(C,b,T,\ell)=\frac{T}{\ell}\left(\frac{L^{2\ell}_{\lambda}}{\sum_{i=1}^{\ell}L^{2i}_{\lambda}}\right)\left(\frac{\lambda C}{b}\right)^{2},$
for some finite $L_{\lambda}\leq\sqrt{3}$ and small enough $\lambda$. Furthermore, we show that the bound obtained in the third regime converges to previously established bounds in the convex setting when each $f_{i}$ becomes increasingly more convex (specifically when $m\to 0$ in contribution (iii)).
Lipschitz continuity. It is worth re-emphasizing that we do not require each $f_{i}$ in (1) to be Lipschitz continuous in order to bound $\nabla f_{i}$ (see, for example, [8, 2, 7] which do require this assumption). As a consequence, our analysis is applicable to a substantially wider class of non-Lipschitzian objective functions, including (a) global or piecewise quadratic functions on $\mathbb{R}^{n}$, (b) the function $f(x)=x\log x$ on $1,\infty)$ and its multivariate extensions, and (c) double integrals of bounded sigmoidal functions, e.g., $f(x)=\int\log\cosh x\,dx=\int\int\tanh x\,dx$, and their multivariate extensions.
1.1 Mathematical techniques
Weak convexity. We generalize existing analyses in the convex and twice-differentiable setting (see, for example, [8, 7]) to the weakly-convex (and possibly nonsmooth) composite setting. Crucially, our approach avoids assuming the existence of the Hessian for the smooth part of the objective function (see, for example, Lemma 3.2 and Proposition 3.4). This generalization is achieved by leveraging recently developed optimal transport techniques [8] and a curvature-based characterization of nonconvex functions described in Section 3. This characterization specifically enables us to show how our bounds converge to similar ones in the convex setting by appropriately scaling a weak convexity parameter.
Upper curvature. To facilitate a more general analysis, we assume that there exists a constant $M>0$ such that each $f_{i}+M\|\cdot\|^{2}/2$ is convex (see Section 3 for an in-depth discussion). This assumption, together with weak-convexity, is mild in that it is implied by assuming each $\nabla f_{i}$ is $L$-Lipschitz continuous for a large enough $L$. On the other hand, it is essential for establishing that DP-SGD is a Lipschitz operator for small enough stepsize (see Proposition 3.4) and is classical in the analysis of first-order optimization algorithms.
Proximal operator. To analyze the effect of gradient clipping, we utilize important properties about proximal operators, such as strict non-expansiveness and uniqueness for closed convex functions. These properties are specifically applied when we view both the projection and clipping steps of DP-SGD as instances of a particular proximal evaluation. Conversely, our analysis may be applied to more general forms of DP-SGD where the gradient clipping step is replaced by any sequence of proximal evaluations (see, for example, Proposition 3.4).
Shifted Rényi divergence. We generalize the approach of shifted Rényi divergences in [8, 2] from nonexpansive [^3] In [8, 2], the authors refer to the relevant operators as contractions even though they are only $1$-Lipschitz, i.e., nonexpansive. operators to $L$-Lipschitz operators for $L>1$. More precisely, while we still sum a sequence of parameterized shifted Rényi divergence bounds, we introduce more parameters compared to previous works to deal with the issue of nonconvexity. Furthermore, we show that in the setting of one dataset pass, the optimal choice of parameters is obtained by solving a related quadratic programming problem (see Appendix B). For the case of multiple dataset passes, we use the optimal parameter in the single dataset setting together with some nice properties of cyclic DP-SGD batch updates (see Proposition 2.2).
1.2 Related works
Recent approaches [8, 7, 6, 2] directly bound the Rényi divergence of the last iterate, but these approaches make strong assumptions that limit their practical viability. In particular, papers [8, 7] only consider the convex setting. Furthermore, they assume Lispchitz continuity of the loss function, which consequently bounds the loss gradients. While paper [6] studies multiple epochs or passes over the data, their results only apply to the smooth and strongly convex setting. Similarly, results in paper[2] consider only convex Lipschitz smooth functions.
Paper [3] analyzes the privacy loss dynamics for nonconvex functions, but their analysis differs from ours in two ways. First, they assume that their DP-SGD batches are obtained by Poisson sampling or sampling without replacement. Second, their results require numerically solving a minimization problem that can be hard for certain parameters and thus their bound is unclear to converge. Tangentially, prior works on DP-SGD accounting often rely on loose bounds that assume the release of intermediate updates [4, 1]. These works rely in differential privacy advanced composition results [9], resulting in noise with standard deviation that scales as $\sqrt{T}$ [4, 1].
In Table 1.2, we compare our bounds against those established using the “Privacy Amplification by Iteration” (PABI) argument in [8] for multi-epoch noisy-SGD under similar smoothness and scaling assumptions. Note that the multi-epoch noisy-SGD algorithm in [8, Algorithm] only considers the case of $E=\ell$ and does not consider batched gradients. As a consequence of the latter, its corresponding RDP upper bound does not depend on the batch size $b$.
Convex Nonconvex $(d_{h}=\infty)$ Nonconvex $(d_{h}<\infty)$ Assumptions
PABI[8] $O\left(\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda Q}{\sigma}\right]^{2}\right)$ - - $f$ is $Q$-Lipschitz; no clipping; no batching
This work $O\left(\dfrac{\alpha E}{\ell}\left[\dfrac{\lambda C}{b\sigma}\right]^{2}\right)$ $O\left(\dfrac{\alpha EL_{\lambda}^{2\ell}}{\sum_{i=0}^{\ell-1}L_{\lambda}^{2i}}\left[\dfrac{\lambda C}{b\sigma}\right]^{2}\right)$ $O\left(\dfrac{\alpha}{\sigma^{2}}\left[L_{\lambda}d_{h}+\dfrac{\lambda C}{b}\right]^{2}\right)$ clipping with an $\ell_{2}$ bound of $C$; batch size $b$
Table 1: Asymptotic $\alpha$-Rényi differential privacy upper bounds of the last iterate after $T$ iterations of DP-SGD, when $f_{i}$ is Lipschitz smooth and weakly-convex for $i\in[k]$, $\lambda$ is the stepsize, $L_{\lambda}=\Theta(\sqrt{1+\lambda m})$ for a weak-convexity parameter $m$ (see (9)), $\sigma$ is the standard deviation of the Gaussian noise, $\ell$ is the number of iterations in one dataset pass, $E$ is the number of dataset passes, and $d_{h}$ is the diameter of the domain of $h$.
1.3 Organization
Section 2 presents some important properties about a specific family of randomized operators. Section 3 states some additional assumptions on (1) and presents some crucial properties for an SGD-like update. Section 4 presents the main RDP bounds of this work. Finally, the appendices at the end of this paper contain technical proofs and extended discussions about the main results.
1.4 Notation
@@
Let $n]=\{1,\ldots,n\}$ for any positive integer $n$. Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{R}^{n}=\mathbb{R}\times\cdots\times\mathbb{R}$ denote the $n$-fold Cartesian product of $\mathbb{R}$. Let $(\langle\cdot,\cdot\rangle,\mathbb{R}^{n})$ denote a Euclidean space over $\mathbb{R}^{n}$ and denote $\|\cdot\|:=\sqrt{\langle\cdot,\cdot\rangle}$ to be the induced norm. The domain of a function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is $\operatorname*{dom}\phi:=\{z\in\mathbb{R}^{n}:\phi(z)<\infty\}$. The function $\phi$ is said to be proper if $\operatorname*{dom}\phi\neq\emptyset$. A function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is said to be lower semicontinuous if $\liminf_{x\to x_{0}}\phi(x)\geq\phi(x_{0})$. The set of proper, lower semicontinuous, convex functions over $\mathbb{R}^{n}$ is denoted by $\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. An operator ${\cal T}:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ is said to be $L$-Lipschitz if $\|{\cal T}(x)-{\cal T}(y)\|\leq L\|x-y\|$ for all $x,y\in\mathbb{R}^{n}$, and $\cal T$} is said to be nonexpansive if it is 1-Lipschitz. The proximal operator ${\rm prox}_{\psi}:\mathbb{R}^{n}\mapsto\mathbb{R}^{n}$ for $\psi\in\overline{{\rm Conv}}\,(\mathbb{R}^{n})$ is defined as
@@
$\operatorname{prox}_{\psi}(z_{0})=\operatorname*{argmin}_{z\in\mathbb{R}^{n}}\left\{\psi(z)+\frac{1}{2}\|z-z_{0}\|^{2}\right\}\quad\forall z_{0}\in\mathbb{R}^{n},$
@@
and it is well-known that ${\rm prox}_{\psi}(\cdot)$ is nonexpansive (see, for example [5, Theorem 6.42]).
@@
$D_{\alpha}(\mu\|\nu)=\frac{1}{\alpha-1}\log\int\left[\frac{\mu(x)}{\nu(x)}\right]^{\alpha}\nu(x)\,dx$

### v3

# Privacy of the last iterate in cyclically-traversed DP-SGD on nonconvex composite losses

@@ Abstract
Differentially-private stochastic gradient descent (DP-SGD) is a family of iterative machine learning training algorithms that privatize gradients to generate a sequence of differentially-private (DP) model parameters. It is also the standard tool used to train DP models in practice, even though most users are only interested in protecting the privacy of the final model. Tight DP accounting for the last iterate would minimize the amount of noise required while maintaining the same privacy guarantee and potentially increasing model utility. However, last-iterate accounting is challenging, and existing works require strong assumptions not satisfied by most implementations. These include assuming (i) the global sensitivity constant is known — to avoid gradient clipping; (ii) the loss function is Lipschitz or convex; and (iii) input batches are sampled randomly.
In this work, we forego any unrealistic assumptions and provide privacy bounds for the most commonly used variant of DP-SGD, in which data is traversed cyclically, gradients are clipped, and only the last model is released. More specifically, we establish new Rényi differential privacy (RDP) upper bounds for the last iterate under realistic assumptions of small stepsize and Lipschitz smoothness of the loss function. Our general bounds also recover the special-case convex bounds when the weak-convexity parameter of the objective function approaches zero and no clipping is performed. The approach itself leverages optimal transport techniques for last iterate bounds, which is a nontrivial task when the data is traversed cyclically and the loss function is nonconvex.
@@ 1 Introduction
Differential privacy (DP) is an approach to capture the sensitivity of an algorithm to any individual user’s data and is frequently used in both industrial and government applications (see the book by [15] for a rich introduction). Given a possibly nonprivate computation $f$, a desired level of DP (or privacy budget) $\varepsilon$ is generally achieved by bounding the global sensitivity [^3] The maximum change in the mechanism’s output caused by changes in a single user or data point. of $f$ and then adding noise to its output. This noise is typically calibrated to the sensitivity and $\varepsilon$ in order to obscure the contributions of a single input example. Conversely, given a mechanism $\cal A$ for making a computation differentially private, a method for determining the level of DP obtained by $\cal A$ is often called a DP accounting method.
Differentially-private stochastic gradient descent (DP-SGD) refers to a family of popular first-order methods for training model weights with DP[10, 27, 6, 1]. At a high level, a DP-SGD method first computes the gradients of a given set of per-example loss functions with respect to the model weights and applies an algorithm $\cal A$} to obtain a private gradient $\cal G$. The private gradient $\cal G$ is then used in some first-order optimization scheme, e.g., SGD, Adam, or AdaGrad, to update model weights. More precisely, $\cal A$} consists of (i) scaling the per-example loss gradients (a.k.a. gradient clipping) to reduce sensitivity, (ii) adding independent and identically distributed (i.i.d.) Gaussian noise $\cal Z$ to each of the scaled gradients, and (iii) summing the noised gradients to obtain $\cal G$. In general, the higher the variance of $\cal Z$ is, the lower the utility of the final trained model.
Depending on the optimization scheme, and the assumptions on how the user-level loss functions are obtained, existing DP accounting methods for DP-SGD can differ significantly. For example, when only the last iterate of DP-SGD is released, existing accounting methods require both sophisticated machinery and numerous strong assumptions to provide tight DP bounds. Some of these strong assumptions, that almost never hold in practice, include (i) the input data is sampled randomly at each DP-SGD iteration, (ii) the loss functions are convex, (iii) the global DP-SGD sensitivity is known beforehand, and (iv) the intermediate model weights are bounded.
This work develops tighter privacy analyses for last-iterate DP-SGD under more realistic settings than existing works. Consequently, our analyses enable implementations of DP-SGD that apply Gaussian noise $\cal Z$ with lower variance than existing work, and as a consequence, obtain higher utility at the same privacy budget $\varepsilon$. More specifically, we develop a family of Rényi DP (RDP) bounds on the last iterate of DP-SGD, which are novel in that they:
@@
    do not assume knowledge of the global sensitivity constant and, hence, are valid with or without gradient clipping;
@@
    hold for both the nonconvex and convex settings under significantly fewer assumptions than other works;
@@
    are parameterized by a weak convexity [^4] A function $f$ is $m$-weakly-convex if $f+m\|\cdot\|^{2}/2$ is convex. parameter $m\geq 0$, for which one of the bounds smoothly converges to a similar one in the convex setting as $m\to 0$.
1.1 Background
We begin by formally stating the problem of interest, describing common terminology and notation, and specifying the DP-SGD variant under consideration. We then briefly describe the Privacy Amplification by Iteration (PABI) argument of [17] and discuss the difficulties of generalizing this argument to more practical settings.
Problem of interest. We develop RDP bounds for the last iterate of a DP-SGD variant applied to the nonsmooth composite optimization problem
@@ 1.1 Background
where $h$ is convex and proper lower-semicontinuous and $f_{i}$ is continuously differentiable on the domain of $h$. Notice that the assumption on $h$ encapsulates (i) common nonsmooth regularization functions such as the $\ell_{1}$-norm $\|\cdot\|_{1}$, nuclear matrix norm $\|\cdot\|_{*}$, elastic net regularizer and (ii) indicator functions on possibly unbounded closed convex sets. A common setting in practice is when $(1/k)\sum_{i=1}^{k}f_{i}(x)$ corresponds to a softmax cross-entropy loss function and $h(x)$ corresponds to an $\ell_{1}$- or $\ell_{2}$-regularization function.
Common terminology. An input data collection $X=\{x_{i}\}_{i=1}^{k}$ is said to be traversed cyclically (or cyclically-traversed) in batches $\{B_{t}\}$ of size $b$ if $B_{t}$ contains $\{x_{b(t-1)+1},\ldots,x_{bt}\}$ for the first $t\leq k/b$ batches [^5] For simplicity, we assume $b$ divides $k$ throughout., and the the rest of the batches cycle between $B_{1},\ldots,B_{k/b}$ in order. Cyclically-traversed DP-SGD is a variant of DP-SGD where the input data is traversed cyclically [^6] Cyclically traversed is also known in the literature as incremental gradient [23].. A dataset pass occurs when the input data (e.g., $X$ above) in a cyclically-traversed DP-SGD run has been used, and the next batch of inputs is the same as the first batch of inputs at the beginning of the dataset pass. Gradient clipping is the process of orthogonally projecting a gradient vector in $\mathbb{R}^{n}$ to a Euclidean ball of radius $C$ centered at the origin. The parameter $C$ is typically called the $\ell_{2}$-norm clip value. In this work, we say a function is a randomized operator if it consists of applying some deterministic operator to an input and adding random noise to resulting output. An operator $\cal T$} is said to be $L$-Lipschitz if $\|{\cal T}(x)-{\cal T}(y)\|\leq L\|x-y\|$ for every $x$ and $y$, and $\cal T$} is said to be nonexpansive if it is 1-Lipschitz.
@@
Common notation. Let $n]=\{1,\ldots,n\}$ for any positive integer $n$. Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{R}^{n}=\mathbb{R}\times\cdots\times\mathbb{R}$ denote the $n$-fold Cartesian product of $\mathbb{R}$. Let $(\langle\cdot,\cdot\rangle,\mathbb{R}^{n})$ denote a Euclidean space over $\mathbb{R}^{n}$ and denote $\|\cdot\|:=\sqrt{\langle\cdot,\cdot\rangle}$ to be the induced norm. The domain of a function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is $\operatorname*{dom}\phi:=\{z\in\mathbb{R}^{n}:\phi(z)<\infty\}$. The function $\phi$ is said to be proper if $\operatorname*{dom}\phi\neq\emptyset$. A function $\phi:\mathbb{R}^{n}\mapsto(-\infty,\infty]$ is said to be lower semicontinuous if $\liminf_{x\to x_{0}}\phi(x)\geq\phi(x_{0})$. The set of proper, lower semicontinuous, convex functions over $\mathbb{R}^{n}$ is denoted by $\overline{{\rm Conv}}\,(\mathbb{R}^{n})$. The clipping operator is given by
${\rm Clip}_{C}(y):=y\cdot\min\left\{1,\frac{C}{\|y\|}\right\},$ (2)
and the proximal operator for a proper convex function $\psi$ is defined as
@@
$\operatorname{prox}_{\psi}(z_{0})=\operatorname*{argmin}_{z\in\mathbb{R}^{n}}\left\{\psi(z)+\frac{1}{2}\|z-z_{0}\|^{2}\right\}\quad\forall z_{0}\in\mathbb{R}^{n}.$ (3)
@@
It is well-known that ${\rm prox}_{\psi}(\cdot)$ is nonexpansive (see, for example [7, Theorem 6.42]) and that ${\rm Clip}_{C}(y)$ is the proximal operator for the (convex) indicator function of the set $\{x:\|x\|\leq C\}$.
@@
$D_{\alpha}(\mu\|\nu)=\frac{1}{\alpha-1}\log\int\left[\frac{\mu(x)}{\nu(x)}\right]^{\alpha}\nu(x)\,dx,$ (4)